OpenStudy (anonymous):

Angles 7 and 8 are vertical angles, <7, congruent sign <8,m<7 =70 degrees, and m<8 = (3x+10) degrees ? whats the value of x

OpenStudy (texaschic101):

vertical angles are equal. so if angle 7 = 70 degrees, then angle 8 is 70 degrees. 3x + 10 = 70 -- subtract 10 3x = 70 - 10 3x = 60 -- divide by 3 x = 20

OpenStudy (anonymous):

thankyou @texaschic101

OpenStudy (texaschic101):

no problem :)